# Large Language Model Can Continue Evolving From Mistakes

### v3

@@ 3.3 Supplemental Training Data
(1) Supplemental Corpus($C_{\underline{S}}$): Supplemental Corpus obtained from multiple data sources contains background knowledge that the LLM lacks and is used similarly to continuous pre-training Gupta et al. (2023).
Example 1: A Case of Normative Corpus
@@
Table 1: A case of a Normative Corpus, designed to illustrate its corpus format.
Example 2: A Case of Extractive Corpus
@@
Table 2: A case of a Extractive Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplemental text has been omitted for clarity.
(2) Normative Corpus($C_{\underline{N}}$): To equip the LLM with the ability to respond to multiple-choice questions in a normative format, we sample multiple-choice questions that are independent of the test set and process them into the format shown in Table 1 to build Normative Corpus. Prior to supplemental training, we finetune the LLM using Normative Corpus(i.e. pre-finetuning), facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of Normative Corpus during the supplemental training to prevent the LLM’s response skills from declining.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To improve the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample multiple-choice questions that are independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
(4) Review Corpus($C_{\underline{R}}$): The responses of the LLM tend to be diverse under high temperature Caccia et al. (2018). During repeated QA testing, LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting the accuracy.
Therefore, to strengthen the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly of the LLM in the previous round and process some of them in Normative Corpus format and the rest in Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus can be described as:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round of LLM. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, to preserve the model’s general capabilities on other tasks, we use the data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
unfiltered
as General Corpus.
Supplemental Training Set Construction Strategy. We have developed three innovative strategies for constructing supplemental training sets by combining five types of corpus. The strategies are detailed below.
(1) Plain Strategy: To mitigate the potential catastrophic forgetting and to maintain the LLM’s ability to provide normative responses, we mix three types of corpus to construct the supplemental training set:
$D_{P}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges various types of corpus together and $D_{P}$ denotes the supplemental training set constructed via the Plain Strategy.
(2) Extractive Strategy: To enhance the LLM’s ability to capture and comprehend knowledge from supplemental material, we integrate the Extractive Corpus:
$D_{E}=Concat(C_{S},C_{N},C_{G},C_{E}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ denotes the supplemental training set constructed via the Extractive Strategy.
(3) Review Strategy: To enhance the model’s retention of acquired knowledge and prevent correct responses from being altered to incorrect ones over multiple iterations, we integrate the Review Corpus:
$D_{R}=Concat(C_{S},C_{N},C_{G},C_{E},C_{R}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ denotes the supplemental training set constructed via the Review Strategy.
4 Experimental Setup
4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.
Xiezhi, a comprehensive evaluation suite designed to assess holistic domain knowledge. It comprises multiple-choice questions across 516 diverse disciplines, ranging from 13 different subjects. Gu et al. (2023).
CMMLU contains about 12,000 questions spanning elementary to advanced professional levels across 67 subjectsLi et al. (2024).
In our experiments, we select 27,265 multiple-choice questions from Xiezhi and 11,917 multiple-choice questions from CMMLU. Appendix Table 7 shows the details of the data division.
4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, and Qwen-7B-Chat [^6] https://huggingface.co/Qwen/Qwen-7B-Chat. These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
CuteGPT-13B is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023).
ChatGLM-6B is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022).
Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
4.3 Metrics
The effectiveness of supplemental training will be assessed using the following metrics:
ACC: it represents the accuracy of the answer of LLM.
5 Experiments
We construct various supplemental training sets and their corresponding control groups for the experiments, following the method described in Section 3.3. To differentiate between models trained on different supplemental training sets, we establish the following naming conventions:
INIT: it represents the model, which has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-N: N denotes Normative, representing the model that has undergone supplemental training using the training set constructed by the Plain Strategy.
[…]
We append suffixes to the aforementioned naming conventions to differentiate the data sources of the supplemental corpus. These suffixes include: ‘Wiki’, ‘Bing’ and ‘Mix’ where ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For instance, ‘E_Wiki’ denotes the model supplemented trained using the dataset constructed through the Extractive Strategy. The Supplemental Corpus of this dataset is collected from Wikipedia.
5.1 Main Results
We conduct experiments to analyze the accuracy after applying CEM on the test data. The results are shown in Figure 2. Our observations are summarized as follows:
Figure 2: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplemental training set is constructed using the Plain Strategy.
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two given tasks. This implies that despite the high availability of supplemental data and the small size of the training set, the CEM method effectively complements the model’s knowledge deficiencies. For instance, the ChatGLM-6B model shows an improvement of up to 16.08% on the CMMLU task.
Secondly, the effectiveness of tuning LLMs with CEM depends on their innate abilities and also varies across LLMs with different parameter scales. Among the mentioned open-source models, ChatGLM-6B, which has the fewest parameters, exhibits the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, with the highest initial answer accuracy and the most parameters, shows a smaller improvement of up to 3.63% compared to the smaller LLM.
Construction Strategy Analysis. To investigate the effectiveness of each component in supplemental training set, we conduct experiments on LLMs trained using various dataset construction strategies, the results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies. ’Avg of $R_{\alpha}$’ and ’Best of $R_{\alpha}$’ represent the average and best results, respectively, of the Review Strategy with different $\alpha$.
Firstly, Extractive Corpus and Review Corpus can significantly boost LLMs’ performance. CEM with Extractive Corpus ($C_{E}$) outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., E_Mix shows a 3.11% increase in accuracy to its control group E’_Mix, and a 10.96% increase compared to INIT. Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{\alpha=0,0.5,1}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, R0.5_Wiki shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, R’_Wiki, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the INIT.
The results indicate an improvement in W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the corpus. This enhancement leads to more effective error correction. Additionally, the Review Corpus significantly reduced R2W, indicating suppression of the model’s tendency to forget correct knowledge.
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. The effectiveness of CEM depends on the recall rate when searching corpus and the quality of supplemental corpus, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Figure 3 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplemental training set constructed by Review Strategy, R_Mix has lower accuracy than R_Wiki and R_Bing despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplemental corpus from different sources. Additionally, an excess of supplemental corpus may dilute the proportion of General, Extractive, and Normative Corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of the Review Corpus. According to Table 6 in Appendix, R0, R0.5 and R1 are similar in their ability to enhance LLMs’ performance. This may be because the models themselves have already mastered the knowledge, and replaying it in either format serves to consolidate that knowledge. Therefore, if not specified, R will refer to the R0.5 Strategy in the following.
5.2 Multiple Iterations Analysis
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction strategy as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
We conduct a second round of supplemental training on CuteGPT-13B on Xiezhi task using the same dataset construction strategy and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extractive Strategy and Review Strategy are effective in improving the accuracy of the model. For instance, R_Wiki shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to INIT.
Furthermore, the Review Strategy consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction strategies, e.g., the R2W metric relative to INIT of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Strategy performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
6 Discussion
Why can the Extractive Corpus be effective? We observe that models employing the Extractive Corpus exhibit higher accuracy and W2R metric. Unlike the Review Corpus, where improving model performance is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, whereas the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions may be a structured induction of supplemental material, where the Extractive Corpus explicitly links complex information within the corpus to the resulting induction (i.e., the question and options). Questions and options can act as cues and guides for the model to gather relevant information from the corpus, considering how this knowledge is accessed through questions. This, in turn, can enhance the model’s ability to extract useful information from the corpus while eliminating redundant and irrelevant information.
7 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplemental corpora from multiple data sources, (3) constructing a supplementl training set and fine-tuning the LLM. Furthermore, we introduce multiple strategies for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
8 Limitations
This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplemental training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
Appendix A Process of Collecting Corpora from Multiple Sources

### v4

@@ 3.3 Supplemental Training Data
(1) Supplemental Corpus($C_{\underline{S}}$): The supplemental corpus, obtained from multiple data sources, contains knowledge that the LLM lacks and is used as CPT data.
Below are the three types of CIT data:
(2) Normative Corpus($C_{\underline{N}}$): To enable the LLM to answer multiple-choice questions in a normative format, we sample questions independent of the test set and process them into the format shown in Table 1 to build the Normative Corpus. Prior to supplemental training, we finetune the LLM using the Normative Corpus (i.e., pre-finetuning), which facilitates extracting answers from the LLM’s responses to calculate initial performance and allows for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of the Normative Corpus during supplemental training.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
(4) Review Corpus($C_{\underline{R}}$):
The responses of the LLM tend to be diverse under high temperature settings Caccia et al. (2018). During repeated QA testing, the LLM may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting accuracy.
To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round. We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus is described as follows:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_unfiltered as the General Corpus to preserve the model’s general capabilities on other tasks.
Supplemental Training Set Construction Strategy. We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data. Specifically, we construct three types of supplemental training sets by combining five types of corpora. The following sections detail these construction strategies.
(1) Plain Strategy:
$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
(2) Extractive Strategy:
$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
(3) Review Strategy:
$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
4 Experimental Setup
4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix Table 6 shows the details of the data division and usage.
4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^2] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^3] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^4] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
4.3 Metrics
The effectiveness of supplemental training will be assessed using the following metrics in the main experiments:
ACC: it represents the accuracy of the answer of LLM.
To analyze the potential performance degradation caused by catastrophic forgetting, we also use the following metrics:
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{i}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
5 Experiments
We construct supplemental training sets and control groups for the experiments, following the strategies described in Section 3.3. To distinguish between models trained on different supplemental training sets, we use the following naming conventions:
INIT represents the model that has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-P: P denotes Plain, representing the model trained with the supplemental training set constructed by the Plain Strategy.
[…]
We append suffixes to indicate the data sources of the Supplemental Corpus: ‘Wiki’, ‘Bing’ and ‘Mix’. ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For example, ‘E-Wiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
5.1 Main Results
We used the Plain Strategy to construct supplementary training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two tasks. This indicates that the CEM method effectively addresses the model’s knowledge deficiencies, even with the small training set and readily available supplemental data. For instance, the ChatGLM-6B improves by up to 16.08% on the CMMLU task.
Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs. For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%. In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
Construction Strategy Analysis. To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies. We set the volume of both Extractive Corpus and Review Corpus to 1,000. The results are shown in Figure 4. Below is a summary of our observations:
[…]
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric. The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. Results in Figure 4 show no significant difference in CEM effectiveness whether using Wikipedia, Bing, or a combination of both. Indeed, R-Mix has lower accuracy than R-Wiki and R-Bing despite having more data. This may be due to redundancy from overlapping sources. Additionally, an excess of supplemental corpus can dilute the proportion of CIT data, reducing their effectiveness in mitigating catastrophic forgetting and enhancing knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of Review Corpus. Table 5 in Appendix shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.
5.2 Multiple Iterations Analysis
We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, using the same training set construction strategy and data source as the first round. We select CIT data independent of the first round but maintain the same data volume. The experimental results presented in Appendix Table 8 show that after the second round of CEM, the accuracy of the LLM’s answers significantly improved across all dataset construction strategies. In the best case, R-Wiki shows a 4.75% improvement in accuracy compared to its previous round and a 17.00% improvement compared to INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of R-Wiki decreases by 7.02% compared to P-Wiki and by 5.90% compared to E-Wiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplementary training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM combined with the random replay method. Here, ER indicates the model’s enhancement rate on the Xiezhi task, while AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
Catastrophic Forgetting Analysis
To analyze catastrophic forgetting that may arise from iterations, we use CuteGPT-13B and test it on Xiezhi, HotpotQA, and GSM8K tasks to evaluate the model’s domain knowledge, reasoning, and mathematical abilities. We employ the CEM method to enhance performance on Xiezhi and assess any performance declines on the other two tasks. Initially, we sample 2,000 instances from each of the three datasets for pre-finetuning, then test to establish baseline. We use a method of random replay to mitigate catastrophic forgetting. Specifically, in each training round, we randomly replay 500 instances from both HotpotQA and GSM8k to replace the original General Corpus.
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi still exhibits significant improvement and maintains an upward trend. Additionally, the combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70% in most cases. However, we observe that CEM-R, with its larger data capacity, achieves significant performance improvements but also shows the greatest decline on other tasks. This may be attributed to the increased volume of CPT data diluting the effect of the replay data. Combining CEM with other advanced replay-based CL methods warrants further investigation.
6 Discussion
Why can the Extractive Corpus be effective? LLMs using the Extractive Corpus show higher accuracy and W2R metrics. Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options). Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions. This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations. In the best case, accuracy enhancement reaches up to 17.00%.
Appendix A Format of CIT Data
Table 1 and Table 2 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
An Example of Normative Corpus
@@ Appendix A Format of CIT Data
Table 1: An example illustrating the format of the Normative Corpus.
An Example of Extractive Corpus
@@
Table 2: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix B Process of Collecting Corpora from Multiple Sources